OpenStudy (anonymous):

Use substitution method to solve the system of equations. X-2y+z=0 2x-3y-4z=-9 X+2y-5z=0

OpenStudy (anonymous):

x+ 2y-5z = 0 -x +2y -z = 0 4y -6z = 0 2x -3y -4z = -9 -2x - 4y +10z = 0 6z = -9 + 7y -3y + 9 = 0 y = -3 6z = -30 z = -5 x = -2y + 5z x = 6 -25 x = -19 ( -19, -3, -5)

hero (hero):

I'm guessing you didn't come here for any math help.

OpenStudy (anonymous):

not truee..I am about to post my question

hero (hero):

I'm just teasing you

OpenStudy (anonymous):

lol